Draw a pie chart for the following data:
Item of expenditure Food Health Clothing Education Savings
Amount spent (in rupees) 3750 1875 1875 1200 7500

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Pie chart is drawn in a circle. So, first, we convert each item of expenditure into angle assuming the total angle in a circle 3600 corresponds to income. Conversion of angle corresponding to the item is calculated by using the formula angle=amount spent on that itemtotal income×3600 After finding the angle for each item of expenditure we divide the circle corresponds to that angle and gives a name to each part. Complete step-by-step solution: Let us assumeI be the income. Let us find the total income by adding the amount spent on each item as ⇒I=3,750+1,875+1,875+1,200+7,500⇒I=16,200 So, the total income is 16,200 rupees. Now, let us calculate the angle for each item of expenditure by using the formula angle=amount spent on that itemtotal income×3600 (i) Let us calculate angle for food By using the above formula we get ⇒angle of food=3,75016,200×3600⇒angle of food=83.34≃830 (ii) Let us calculate angle for health By using the above formula we get ⇒angle of health=1,87516,200×3600⇒angle of clothing=41.67≃420 (iii) Let us calculate angle for clothing By using the above formula we get ⇒angle of clothing=1,87516,200×3600⇒angle of clothing=41.67≃420 (iv) Let us calculate angle for education By using the above formula we get ⇒angle of education=1,20016,200×3600⇒angle of education=26.67≃270 (v) Let us calculate angle for savings By using the above formula we get ⇒angle of savings=7,50016,200×3600⇒angle of savings=166.34≃1660 Now, by making these angles as parts in a circle we get a pie chart as Note: Students may do mistake in following points (i) We need to mention each part with its name. (ii) Give each part to a certain design to avoid confusion (iii) Angle should not be mentioned in the pie diagram but you can mention the amount spent. Above mentioned points need to be taken care of while drawing the pie chart.